### v2

# Bollobás-Erdős-Tuza conjecture for graphs with no induced $K_{s,t}$

Abstract
A widely open conjecture proposed by Bollobás, Erdős, and Tuza in the early 1990s states that for any $n$-vertex graph $G$, if the independence number $\alpha(G)=\Omega(n)$, then there is a subset $T\subseteq V(G)$ with $|T|=o(n)$ such that $T$ intersects all maximum independent sets of $G$. In this paper, we prove that this conjecture holds for graphs that do not contain an induced $K_{s,t}$ for fixed $t\geq s$. Our proof leverages the probabilistic method at an appropriate juncture.
1 Introduction
For a graph $G=(V,E)$, an independent set is a set of vertices in $G$ such that no pair of vertices is adjacent. An independent set is a maximum independent set if it has the largest possible size in the graph $G$, and this size is called the independence number of $G$. Usually we write $\alpha(G)$ as the independence number of $G$. A subset $T\subseteq V(G)$ is a hitting set of the graph $G$ if $T\cap I\neq\emptyset$ holds for each maximum independent set $I\subseteq V(G)$. Define $h(G)$ to be the smallest size of a hitting set $T\subseteq V(G)$. In this paper, we always assume that $n$ is sufficiently large whenever necessary. Bollobás, Erdős and Tuza (see, [6, 7]) proposed the following conjecture.
Conjecture 1 ([6, 7]).
Let $G$ be an $n$-vertex graph with $\alpha(G)=\Omega(n)$, then $h(G)=o(n)$.
According to Hajnal’s observation [14], this conjecture is trivial when $\alpha(G)>\frac{n}{2}$, but becomes challenging in general, even when $\alpha(G)=(\frac{1}{2}-\varepsilon)n$ for any $\varepsilon>0$. A feasible task is to investigate whether this conjecture holds for certain special classes of graphs. For example, Alon [2, 3] demonstrated that every $n$-vertex regular graph $G$ with $\alpha(G)>(\frac{1}{4}+\varepsilon)n$ for any $\varepsilon>0$ satisfies $h(G)\leq O_{\varepsilon}(\sqrt{n\log{n}})$. Recently, Cheng and Xu [5] proved that this conjecture holds for several classes of graphs, including even-hole-free graphs, disk graphs, and circle graphs. We also recommend interested readers to explore some related results and approaches [1, 13].
It has been proven in [5] that for any $n$-vertex even-hole-free graph $G$ with linear-sized independence number, $h(G)\leq\frac{5n}{\log{n}}$. Note that an even-hole-free graph does not contain any induced even cycles of length at least four. It is natural to ask whether we can strengthen the above result by only forbidding a single induced even cycle, such as an induced $4$-cycle. We answer this question for induced $C_{4}$-free graphs, albeit with a slightly weaker upper bound on $h(G)$ compared to the case of even-hole-free graphs. More generally, our main result shows that 1 holds for graphs without induced complete bipartite graph $K_{s,t}$ for any fixed $t\geq s$.
Theorem 1.1.
For fixed positive integers $t\geq s$, let $G$ be an $n$-vertex induced $K_{s,t}$-free graph with $\alpha(G)=\Omega(n)$, then $h(G)=o(n)$.
2 Proof of Theorem 1.1
Let $G$ be an $n$-vertex induced $K_{s,t}$-free graph with $\alpha(G)=cn$ for some $0<c\leq\frac{1}{2}$. Let $\delta>0$ be an arbitrary small number. It suffices to show that $h(G)<\delta n$ holds for any sufficiently large $n$. For a subset $W\subseteq V(G)$ and a vertex $v\in V(G)$, we use $N_{W}(v)$ to denote the neighborhood of $v$ in $W$ and sometimes we also use $N(v)$ if the subscript is clear. For subsets $A,B\subseteq V(G)$, we define $E(A,B)$ as the set of edges with one endpoint in $A$ and the other in $B$.
Since for every vertex $v\in V(G)$, $\{v\}\cup N_{G}(v)$ is a hitting set, we can assume that the minimum degree of $G$ is at least $\delta n-1$. We then arbitrarily pick a maximum independent set $I$. Firstly, we focus on the set of vertices in $V(G)\setminus I$ whose number of neighbors falls within a specific range, more precisely, for $1\leq j\leq\frac{2}{\delta}$, let
$S_{j}:=\{v\in V(G)\setminus I:n(\log n)^{-(10s)^{2j+1}}\leq|N_{I}(v)|<n(\log n)^{-(10s)^{2j-1}}\}.$
Since $|V(G)\setminus I|=(1-c)n$, by pigeonhole principle, there exists some $1\leq j\leq\frac{2}{\delta}$ such that $|S_{j}|\leq\frac{(1-c)\delta n}{2}<\frac{\delta n}{2}$. Fix this $j$.
We pick $k_{j}:=(\log{n})^{(10s)^{2j}}$ vertices in $I$ uniformly at random and let $I_{j}=\{u_{1},u_{2},\ldots,u_{k_{j}}\}\subseteq I$ be the subset consisting of those chosen vertices. For each $s$-tuple $(p_{1},\ldots,p_{s})\in\binom{[k_{j}]}{s}$, let $K^{(p_{1},\ldots,p_{s})}$ be the subgraph induced by the common neighborhood of $u_{p_{1}},\ldots,u_{p_{s}}$. Note that $G$ is induced $K_{s,t}$-free, therefore, for each $s$-tuple $(p_{1},\ldots,p_{s})\in\binom{[k_{j}]}{s}$, we have $\alpha(K^{(p_{1},\ldots,p_{s})})\leq t-1$. Here we remark that the vertices in distinct $K^{(p_{1},\ldots,p_{s})}$ might be overlap, even identical. Let $K\subseteq V(G)$ be a subset consisting all vertices in $K^{(p_{1},\ldots,p_{s})}$ for all $s$-tuples $(p_{1},\ldots,p_{s})\in\binom{[k_{j}]}{s}$.
We then consider the edges between $I$ and $V(G)\setminus(I\cup K\cup S_{j})$, that is, the edge set $E(I,V(G)\setminus(I\cup K\cup S_{j}))$. Notice that this edge set depends on $K$, which in turn depends on the random choice of $I_{j}$. For convenience, let $e:=|E(I,V(G)\setminus(I\cup K\cup S_{j}))|$ be a random variable.
Note that for a vertex $v\in I$, the average number of $|N_{V(G)\setminus(I\cup K\cup S_{j})}(v)|$ is $\frac{e}{cn}$. Therefore, there exists a subset $H\subseteq I$ of size $|H|=(t-1)\binom{k_{j}}{s}+1$ such that $|E(H,V(G)\setminus(I\cup K\cup S_{j})|\leq\frac{\big{(}(t-1)\binom{k_{j}}{s}+1\big{)}e}{cn}$. Fix this subset $H$ and let
$N(H):=\bigcup\limits_{v\in H}N_{V(G)\setminus(I\cup K\cup S_{j})}(v).$
Claim 2.1.
$H\cup N(H)\cup S_{j}$ hits all maximum independent sets of $G$.
Proof of claim.
Suppose there is a maximum independent set $I^{\prime}\subseteq V(G)$ such that $I^{\prime}\cap(H\cup N(H)\cup S_{j})=\emptyset$. Note that $|I^{\prime}\setminus K|\geq cn-(t-1)\binom{k_{j}}{s}$ since a maximum independent set can intersect each $K^{(p_{1},\ldots,p_{s})}$ with at most $t-1$ vertices. Also note that $I^{\prime}\cap N(H)=\emptyset$, therefore $H\cup(I^{\prime}\setminus K)$ is an independent set of size $cn-(t-1)\binom{k_{j}}{s}+(t-1)\binom{k_{j}}{s}+1>cn$, a contradiction to that $I^{\prime}$ is a maximum independent set. ∎
By choices of $H$ and $S_{j}$, we can see that $|H\cup N(H)\cup S_{j}|<(t-1)\binom{k_{j}}{s}+1+\frac{\big{(}(t-1)\binom{k_{j}}{s}+1\big{)}e}{cn}+\frac{\delta n}{2}$. By 2.1, if $(t-1)\binom{k_{j}}{s}+1+\frac{\big{(}(t-1)\binom{k_{j}}{s}+1\big{)}e}{cn}+\frac{\delta n}{2}<\delta n$, that is, if $e<\frac{\delta cn^{2}}{2(t-1)\binom{k_{j}}{s}+2}-cn$, then we are done.
It suffices to provide a desired upper bound on $\mathbb{E}[e]$, which is the expected number of edges in $E(I,V(G)\setminus(I\cup K\cup S_{j}))$. To achieve this, we partition the vertices in $V(G)\setminus(I\cup K\cup S_{j})$ into two groups based on $|N_{I}(v)|$. More precisely, let
$A_{L}:=\{v\in V(G)\setminus(I\cup K\cup S_{j}):|N_{I}(v)|\geq n(\log{n})^{-(10s)^{2j-1}}\},$
and
$A_{S}:=\{v\in V(G)\setminus(I\cup K\cup S_{j}):|N_{I}(v)|<n(\log{n})^{-(10s)^{2j+1}}\}.$
Since $|V(G)\setminus(I\cup K\cup S_{j})|\leq(1-c)n$, the expected number of edges in $E(I,A_{S})$ is less than
$n(\log{n})^{-(10s)^{2j+1}}\cdot(1-c)n<\frac{1}{10}\cdot\bigg{(}\frac{\delta cn^{2}}{2(t-1)\binom{k_{j}}{s}+2}-cn\bigg{)}.$
It remains to estimate the expected number of edges in $E(I,A_{L})$. The key observation is that, for a vertex $v\in V(G)\setminus I$ such that $|N_{I}(v)|\geq n(\log{n})^{-(10s)^{2j-1}}$, the probability that $v\in A_{L}$ is equal to the probability that $I_{j}$ contains at most $s-1$ vertices in $N_{I}(v)$. Therefore, the expected number of edges in $E(I,A_{L})$ is at most
$\begin{split}&\sum\limits_{\begin{subarray}{c}v\in V(G)\setminus I:\\ |N_{I}(v)|\geq n(\log{n})^{-(10s)^{2j-1}}\end{subarray}}\sum\limits_{x=0}^{s-1}\mathbb{P}\bigg{[}|I_{j}\cap N_{I}(v)|=x\bigg{]}\cdot|N_{I}(v)|\\ &\leq\sum\limits_{\begin{subarray}{c}v\in V(G)\setminus I:\\ |N_{I}(v)|\geq n(\log{n})^{-(10s)^{2j-1}}\end{subarray}}\sum\limits_{x=0}^{s-1}\binom{k_{j}}{x}\bigg{(}1-\frac{|N_{I}(v)|}{cn}\bigg{)}^{k_{j}-x}\bigg{(}\frac{|N_{I}(v)|}{cn}\bigg{)}^{x}\cdot cn\\ &\leq c(1-c)n^{2}\cdot\sum\limits_{x=0}^{s-1}\binom{k_{j}}{x}\bigg{(}1-\frac{n(\log{n})^{-(10s)^{2j-1}}}{cn}\bigg{)}^{k_{j}-x}\bigg{(}\frac{cn}{cn}\bigg{)}^{x}\\ &\leq c(1-c)n^{2}\cdot\sum\limits_{x=0}^{s-1}k_{j}^{x}\bigg{(}1-\frac{(\log{n})^{-(10s)^{2j-1}}}{c}\bigg{)}^{k_{j}-x}\\ &\leq\frac{1}{10}\cdot\bigg{(}\frac{\delta cn^{2}}{2(t-1)\binom{k_{j}}{s}+2}-cn\bigg{)}.\end{split}$
By the linearity of expectation, the expected number of $e$ is smaller than $\frac{\delta cn^{2}}{2(t-1)\binom{k_{j}}{s}+2}-cn$, which implies the existence of a choice of $I_{j}$ such that $|H\cup N(H)\cup S_{j}|<\delta n$. The proof is then completed by 2.1.
3 Some remarks
3.1 More on the proof
We originally thought that the following result would be helpful for our proof, as it was proved by Gyárfás, Hubenko, and Solymosi [11] and more recently by Holmsen [16] via a different approach. We also recommend interested readers to refer to [12, 17, 18]. Let $\omega(G)$ denote the size of the largest complete subgraph in a graph, generally referred to as the clique number of the graph $G$.
Theorem 3.1 ([11, 16]).
Let $G$ be an $n$-vertex induced $C_{4}$-free graph with $e(G)\geq\alpha\binom{n}{2}$ for some $\alpha>0$, then there is some $\beta>0$ such that $\omega(G)\geq\beta n$.
We aimed to systematically remove cliques of linear size from an induced $C_{4}$-free graph, with the goal of eventually reaching a contradiction. However, the quantitative result in Theorem 3.1 indicates that repeatedly finding cliques of appropriate sizes is not feasible. Therefore, in the actual proof of Theorem 1.1, we employ the probabilistic method. Nevertheless, we believe that such a result is relevant to our main topic, so we present it here along with a self-contained proof using dependent random choice [10].
Proof of Theorem 3.1.
Let $\beta:=\frac{\alpha^{2}}{128}$. We call $uv$ a missing edge if $uv\notin E(G)$. As $G$ is induced $C_{4}$-free, for any missing edge $uv$, $G[N(u)\cap N(v)]$ must form a clique. We can assume that for every missing edge $uv$, $|N(u)\cap N(v)|<\beta\cdot n$, otherwise we are done. Pick one vertex $x\in V(G)$ uniformly at random, let $U:=N(x)$. It is easy to see that $\mathbb{E}[|U|]=\sum\limits_{v\in V(G)}\frac{|N(v)|}{n}\geq\alpha(n-1)$. Let $Y$ be the random variable counting the number of missing edges in $U$, then we can see the expected number $\mathbb{E}[Y]<\frac{\beta n}{n}(1-\alpha)\binom{n}{2}=\beta(1-\alpha)\binom{n}{2}$, where the first inequality is due to the probability that a missing edge belongs to $U$ is at most $\frac{\beta n}{n}$. Let $Z:=|U|-\frac{\alpha\cdot Y}{\beta(1-\alpha)n}-\frac{\alpha(n-1)}{2}$. By linearity of expectation, $\mathbb{E}[Z]\geq 0$ and therefore there is a choice of $x$ such that $|U|\geq\frac{\alpha(n-1)}{2}$ and $\frac{Y}{|U|}\leq\frac{\beta(1-\alpha)n}{\alpha}$.
Fix this vertex $x$, and take a maximal matching of missing edges in $U=N(x)$, namely, $\tau_{1},\tau_{2},\ldots,\tau_{s}$. Since $G$ is induced $C_{4}$-free, for every pair $\tau_{i}$ and $\tau_{j}$, there is at least one more missing edge sharing one vertex with $\tau_{i}$ and $\tau_{j}$, respectively, which implies that the number of missing edges in $U$ is at least $s+\binom{s}{2}$. One the other hand, we have $s+\binom{s}{2}\leq|Y|\leq\frac{\beta(1-\alpha)n}{\alpha}\cdot|U|$. Also note that $\omega(G)\geq|U|-2s+1$, then the result follows by simple calculations. ∎
Remark 3.2.
Determining the optimal constant dependence between $\alpha$ and $\beta$ in Theorem 3.1 is of interest. The current best-known result is $\beta=(1-\sqrt{1-\alpha})^{2}$, which was shown in [16]
3.2 Graphs without other induced substructures
In general, we can show that 1 holds for those graphs having linear independence number but the number of maximum independent sets is relatively small.
Theorem 3.3.
Let $G$ be an $n$-vertex graph with $\alpha(G)=\Omega(n)$. If the number of distinct maximum independent sets is $2^{o(n)}$, then $h(G)=o(n)$.
There are two simple approaches to showing Theorem 3.3. The first is the same as the proof in [2, Theorem 1.5], one can randomly sample $p:=o(n)$ vertices and the probability that a maximum independent set is not hit is at most $(1-\frac{p}{n})^{cn}$. Using a union bound, then one can show with a positive probability, that the randomly chosen vertices hit all maximum independent sets. The second proof is similar to the proof in [1, Theorem 1.2]. One can build a set system $\mathcal{F}$ consisting of all maximum independent sets in $G$. It then suffices to upper bound the transversal number of $\mathcal{F}$, which in turn is upper bounded by a function of its fractional transversal number and VC-dimension, due to the well-known $\varepsilon$-net theorem in [15]. Furthermore, the fractional transversal number is a constant due to $\alpha(G)=\Omega(n)$, and the VC-dimension is $o(n)$ since the number of maximum independent sets is $2^{o(n)}$. We omit the details and straightforward calculations of both proofs here.
In particular, It was shown in [8, 9] that the number of maximum independent sets in an $n$-vertex graph that does not contain an induced matching of size $t$ is at most $O(n^{2t-2})$. Therefore, for an $n$-vertex graph $G$ with no induced matching of size $t$, if $\alpha(G)=\Omega(n)$, then $h(G)=O(\log{n})$.
We are also interested in whether 1 holds for induced-$H$-free graphs when $H$ is not a complete bipartite graph. For instance, it will be interesting to generalize the result on $4$-cycle to longer even cycles.
Conjecture 2.
For a positive integer $k\geq 3$, let $G$ be an $n$-vertex induced $C_{2k}$-free graph with $\alpha(G)=\Omega(n)$, then $h(G)=o(n)$.
A result of Brandt [4] states that if $G$ is a maximal triangle-free graph without induced $6$-cycle, then every independent set of $G$ is contained in the neighborhood of a vertex. Therefore, the number of maximum independent sets in such a graph is at most $n$, then one can take $O(\log{n})$ vertices to hit all maximum independent sets.